OpenStudy (anonymous):

Find the area of a regular hexagon with side length of 8 m. Round your answer to the nearest tenth.

hero (hero):

There are quite a bit of steps to this

hero (hero):

The formula for area of a polygon is \[A_{polygon} = \frac{Pa}{2}\] Where P = perimeter and a = apothem

OpenStudy (anonymous):

There is a simpler formula, recall that a regular hexagon is made up of 6 equilateral triangle so the area should be \( \large \frac{3 \sqrt{3}}{2}a^2 \) where \( a\) is the length of the side.